# Exact calculation of degrees for lattice equations: a singularity approach

### v2

@@ Acknowledgement
I wish to thank Prof. R. Willox for discussion and comments. This work was supported by a Grant-in-Aid for Scientific Research of Japan Society for the Promotion of Science, JSPS KAKENHI Grant Number 23K12996.
@@ Appendix A Essential choices of domain
In this appendix, we show that there are only four essential choices of domain when one considers the individual degrees with respect to an arbitrary one of the initial variables.
@@
Suppose that a domain $H\subset\mathbb{Z}^{2}$ satisfies the conditions in Definition 3.1 and consider a quad equation (1.1) on $H$. Let $(t_{0},n_{0})\in H_{0}$, $z=x_{t_{0},n_{0}}$ and consider $\deg_{z}x_{t_{0},n_{0}}(z)$, the degree with respect to $z$. Note that by the second condition of Definition 3.1, $(t_{0}+1,n_{0}+1)$ must belong to $H\setminus H_{0}$. Define domains $H^{(1)},H^{(2)},H^{(3)},H^{(4)}\subset\mathbb{Z}^{2}$ as
@@
We check that $x_{s,m}(z)$ ($(s,m)\in H^{(j)}_{0}$), which are the solution on $H$, can be thought of as if they were initial variables (see Figure 9). It is sufficient to show the following two statements:
@@
+++++++++++++++$H^{(2)}$$z$+++++++++
Figure 9. Situation in Theorem A.1. The points marked with “$+$” correspond to the initial boundary of $H^{(2)}$. For example, it follows from this figure that $x_{t_{0}+2,n_{0}}(z)$, which corresponds to a point marked with “$+$” but is not an initial variable in $H$, does not depend on $z=x_{t_{0},n_{0}}$.
@@ Appendix B Substitution and derivatives
In this appendix, let $K$ be an arbitrary field of characteristic 00 and consider the rational function field $K(z,w)$. We allow $\infty$ as the result of substitution throughout this paper. This is why we need some unusual definitions and lemmas about substitution and derivatives.
@@
First, we consider the case $z^{*}\neq\infty$. Let $f(z,w)=\frac{p(z,w)}{q(z,w)}$ where $p,q\in K[z,w]$ are coprime. If $q(z^{*},w)=0$, i.e., $q$ is divisible by $z-z^{*}$, then we set $f(z^{*},w)=\infty$. Note that in this case, $p(z^{*},w)$ cannot become 00. If $q(z^{*},w)\neq 0$, then we define $f(z^{*},w)$ as $f(z^{*},w)=\frac{p(z^{*},w)}{q(z^{*},w)}$.
@@
Since $K$ has characteristic 00, it is easy to show that $\partial_{w}f(z,w)=0$ if and only if $f\in K(z)$.
@@
Therefore, in our settings, it is impossible to completely determine only by $\partial_{w}$ whether a rational function depends on $w$.
@@
Next, we show the converse. Let $\partial_{w}f(z^{*},w)=0$. If $f(z^{*},w)=\infty$, then it is obviously independent of $w$. Therefore, we may assume that $f(z^{*},w)\neq\infty$. In this case, usual properties of the derivative hold and thus $f(z^{*},w)$ does not depend on $w$. If $\partial_{w}\frac{1}{f(z^{*},w)}=0$, one can show in the same way that $g(z^{*},w)$ does not depend on $w$ where $g(z,w)=\frac{1}{f(z,w)}$. ∎
@@ Appendix C Algebraic lemmas
Let $F=F(X;Y_{1},\ldots,Y_{N})\in\mathbb{C}(X;Y_{1},\ldots,Y_{n})$ and let $x^{*}\in\mathbb{C}$. Suppose that the expansion of $F(x^{*}+\epsilon;Y_{1},\ldots,Y_{N})$ at $\epsilon=0$ is
$F(x^{*}+\epsilon;Y_{1},\ldots,Y_{N})=F_{\ell}(Y_{1},\ldots,Y_{N})\epsilon^{\ell}+O\left(\epsilon^{\ell+1}\right),$
@@
Let $F_{\ell}=\frac{p_{\ell}}{q_{\ell}}$ where $p_{\ell},q_{\ell}\in\mathbb{C}[Y_{1},\ldots,Y_{N}]$ are coprime. Since $F_{\ell}$ is not 00 nor $\infty$, neither $p_{\ell}$ nor $q_{\ell}$ is 00 as a polynomial. Since $y_{1}(z^{*}),\ldots,y_{N}(z^{*})$ are algebraically independent over $\mathbb{C}$, we have
@@
Since the statement is trivial in the case of $\operatorname{trdeg}_{\mathbb{C}}\mathbb{C}(y_{1},\ldots,y_{N})=N$, we may assume that the transcendental degree is $N-1$. Let
@@
If $\operatorname{trdeg}_{\mathbb{C}}\mathbb{C}(y_{1},\ldots,y_{N})=N$, then both $F_{1}(y_{1},\ldots,y_{N})$ and $F_{2}(y_{1},\ldots,y_{N})$ are non-zero. Seeking for a contradiction, assume that the transcendental degree is $N-1$ but $F_{1}(y_{1},\ldots,y_{N})=F_{2}(y_{1},\ldots,y_{N})=0$. Let $G$ be the irreducible polynomial in Lemma C.3. Then, $F_{1}$ and $F_{2}$ both belong to $(G)$, which contradicts the coprimeness of $F_{1}$ and $F_{2}$. ∎
@@
Appendix D Divisors on $\mathbb{P}^{1}$
For those who are not familiar with algebraic geometry, we briefly review the basic notion of divisors on $\mathbb{P}^{1}$. In this paper, we use divisors only in order to exactly count the number of preimages with multiplicity. As for the base field $\overline{\mathbb{K}}$, we only use the property that it is algebraically closed. Although the definition of divisors is the same as usual, some notations, such as $\operatorname{div}(f=\alpha\mid T)$, are introduced by the author.
@@ Appendix D Divisors on $\mathbb{P}^{1}$
A divisor on $\mathbb{P}^{1}$ is a finite formal sum over the points of $\mathbb{P}^{1}$ with $\mathbb{Z}$-coefficients. We use $\operatorname{Div}\mathbb{P}^{1}$ to denote the set of divisors on $\mathbb{P}^{1}$.
@@
For a rational function $f\colon\mathbb{P}^{1}\to\mathbb{P}^{1}$, $\alpha\in\mathbb{P}^{1}$ and a subset $T\subset\mathbb{P}^{1}$, we define $\operatorname{div}(f=\alpha\mid T)\in\operatorname{Div}\mathbb{P}^{1}$ as
@@
where $n_{P}$ is the multiplicity of $f=\alpha$ at $P$. Note that the multiplicity is considered to be 00 if $f(P)\neq\alpha$ or $f$ is a constant. We simply write
$\operatorname{div}(f=\alpha)=\operatorname{div}(f=\alpha\mid\mathbb{P}^{1}).$
@@
Let $f\colon\mathbb{P}^{1}\to\mathbb{P}^{1}$ be a rational function and let $(T_{\lambda})_{\lambda\in\Lambda}$ be a family of disjoint subsets of $\mathbb{P}^{1}$. Then, we have
@@
For a divisor $D=\sum_{P\in\mathbb{P}^{1}}n_{P}P$, we define its degree and support as
$\deg D=\sum_{P\in\mathbb{P}^{1}}n_{P},\quad\operatorname{supp}D=\{P\in\mathbb{P}^{1}\mid n_{P}\neq 0\}.$
@@
For a rational function $f\colon\mathbb{P}^{1}\to\mathbb{P}^{1}$ and an arbitrary value $\alpha\in\mathbb{P}^{1}$, the following equality holds:
@@
where the right-hand side is the degree of $f$ as a rational function.

### v3

@@ Acknowledgement
I wish to thank Prof. R. Willox for discussion and comments. I also thank the anonymous referee whose comments helped improve the presentation of the paper. This work was supported by a Grant-in-Aid for Scientific Research of Japan Society for the Promotion of Science, JSPS KAKENHI Grant Number 23K12996.
Data availability
All numerical calculations supplementing the results of this study are included in the article.
@@ Appendix A Essential choices of domain
In this appendix, we show that there are only four essential choices of domain when one considers the degrees with respect to one of the initial variables.
@@
Suppose that a domain $H\subset\mathbb{Z}^{2}$ (see Definition 3.2) and consider a quad equation (1.2) on $H$. Let $(t_{0},n_{0})\in H_{0}$, $z=x_{t_{0},n_{0}}$ and consider $\deg_{z}x_{t,n}(z)$, the degree with respect to $z$. Note that by the second condition of Definition 3.2, $(t_{0}+1,n_{0}+1)$ must belong to $H\setminus H_{0}$. Define domains $H^{(1)},H^{(2)},H^{(3)},H^{(4)}\subset\mathbb{Z}^{2}$ as
@@
[…]
@@
$H^{(2)}$zz++++++
Figure 9. Situation in Theorem A.1. The points marked with “++” correspond to the initial boundary of $H^{(2)}$. For example, it follows from this figure that $x_{t_{0}+2,n_{0}}(z)$, which corresponds to a point marked with “++” but is not an initial variable in $H$, does not depend on $z=x_{t_{0},n_{0}}$.
@@ Appendix B Substitution and derivatives
In this appendix, let $K$ be an arbitrary field of characteristic 0 and consider the rational function field $K(z,w)$. We allow $\infty$ as the result of substitution throughout this paper. This is why we need some unusual definitions and lemmas about substitution and derivatives.
@@
First, we consider the case $z^{*}\neq\infty$. Let $f(z,w)=\frac{p(z,w)}{q(z,w)}$ where $p,q\in K[z,w]$ are coprime. If $q(z^{*},w)=0$, i.e., $q$ is divisible by $z-z^{*}$, then we set $f(z^{*},w)=\infty$. Note that in this case, $p(z^{*},w)$ cannot become 0. If $q(z^{*},w)\neq 0$, then we define $f(z^{*},w)$ as $f(z^{*},w)=\frac{p(z^{*},w)}{q(z^{*},w)}$.
@@
Since $K$ has characteristic 0, it is easy to show that $\partial_{w}f(z,w)=0$ if and only if $f\in K(z)$.
@@
Therefore, in our settings, it is impossible to determine only by $\partial_{w}$ whether a rational function depends on $w$.
@@
Next, we show the converse. Let $\partial_{w}f(z^{*},w)=0$. If $f(z^{*},w)=\infty$, then it is independent of $w$. Therefore, we may assume that $f(z^{*},w)\neq\infty$. In this case, usual properties of the derivative hold and thus $f(z^{*},w)$ does not depend on $w$. If $\partial_{w}\frac{1}{f(z^{*},w)}=0$, one can show in the same way that $g(z^{*},w)$ does not depend on $w$ where $g(z,w)=\frac{1}{f(z,w)}$. ∎
@@ Appendix C Algebraic lemmas
Let $F=F(X;Y_{1},\ldots,Y_{N})\in\mathbb{C}(X;Y_{1},\ldots,Y_{n})$ and let $x^{*}\in\mathbb{C}$. Suppose that the expansion of $F(x^{*}+\varepsilon;Y_{1},\ldots,Y_{N})$ at $\varepsilon=0$ is
$F(x^{*}+\varepsilon;Y_{1},\ldots,Y_{N})=F_{\ell}(Y_{1},\ldots,Y_{N})\varepsilon^{\ell}+O\left(\varepsilon^{\ell+1}\right),$
@@
Let $F_{\ell}=\frac{p_{\ell}}{q_{\ell}}$ where $p_{\ell},q_{\ell}\in\mathbb{C}[Y_{1},\ldots,Y_{N}]$ are coprime. Since $F_{\ell}$ is not 0 or $\infty$, neither $p_{\ell}$ nor $q_{\ell}$ is 0 as a polynomial. Since $y_{1}(z^{*}),\ldots,y_{N}(z^{*})$ are algebraically independent over $\mathbb{C}$, we have
@@
Since the statement is trivial in the case of $\operatorname{trdeg}_{\mathbb{C}}\mathbb{C}(y_{1},\ldots,y_{N})=N$, we may assume that the transcendental degree is N1N-1. Let
@@
If $\operatorname{trdeg}_{\mathbb{C}}\mathbb{C}(y_{1},\ldots,y_{N})=N$, then both $F_{1}(y_{1},\ldots,y_{N})$ and $F_{2}(y_{1},\ldots,y_{N})$ are non-zero. Seeking a contradiction, assume that the transcendental degree is N1N-1 but $F_{1}(y_{1},\ldots,y_{N})=F_{2}(y_{1},\ldots,y_{N})=0$. Let $G$ be the irreducible polynomial in Lemma C.3. Then, $F_{1}$ and $F_{2}$ both belong to $(G)$, which contradicts the coprimeness of $F_{1}$ and $F_{2}$. ∎
@@
Appendix D Divisors on $\mathbb{P}^{1}(K)$
For those who are not familiar with algebraic geometry, we briefly review the basic notion of divisors on $\mathbb{P}^{1}(K)$, where $K$ is an algebraically closed field. In this paper, we use divisors only in order to exactly count the number of preimages with multiplicity. Although the definition of divisors is the same as usual, some notations, such as $\operatorname{div}(f=\alpha\mid T)$, are introduced by the author.
@@ Appendix D Divisors on $\mathbb{P}^{1}(K)$
Let $f\colon\mathbb{P}^{1}(K)\to\mathbb{P}^{1}(K)$ be a rational function, $P\in\mathbb{P}^{1}(K)$ and $\alpha\in\mathbb{P}^{1}(K)$. Let $z$ denote the inhomogeneous coordinate of $\mathbb{P}^{1}(K)$. We define the multiplicity of $f=\alpha$ at $P$, say $n_{P}$, as follows.
  1. (1)
    If $f(P)\neq\alpha$, then we define $n_{P}=0$.
  2. (2)
    If $f$ is the constant function $\alpha$, then we define $n_{P}=0$.
    For the remaining cases, we assume that $f(P)=\alpha$ and $f$ is not a constant function.
  3. (3)
    If neither $P$ nor $\alpha$ is $\infty$, then $f(z)-\alpha$ is expressed as
    $f(z)-\alpha=(z-P)^{m}\frac{g(z)}{h(z)},$
    where $g(P)h(P)\neq 0$ and $m>0$. We define $n_{P}=m$.
  4. (4)
    If $P=\infty$ and $\alpha\neq\infty$, then $f(z)-\alpha$ is expressed as
    $f(z)-\alpha=\frac{g(z)}{h(z)},$
    where $g(z)$ and $h(z)$ are coprime. We define $n_{P}=\deg_{z}h(z)-\deg_{z}g(z)$.
  5. (5)
    If $P\neq\infty$ and $\alpha=\infty$, then $f(z)$ is expressed as
    $f(z)=(z-P)^{-m}\frac{g(z)}{h(z)},$
    where $g(P)h(P)\neq 0$ and $m>0$. We define $n_{P}=m$.
  6. (6)
    If both $P$ and $\alpha$ are $\infty$, then $f(z)$ is expressed as
    $f(z)=\frac{g(z)}{h(z)},$
    where $g(z)$ and $h(z)$ are coprime. We define $n_{P}=\deg_{z}g(z)-\deg_{z}h(z)$.
The reason we define $n_{P}=0$, rather than $\infty$, for the constant function $f=\alpha$ is that this choice ensures Proposition D.6.
@@
A divisor on $\mathbb{P}^{1}(K)$ is a finite formal sum over the points of $\mathbb{P}^{1}(K)$ with $\mathbb{Z}$-coefficients. We use $\operatorname{Div}\mathbb{P}^{1}(K)$ to denote the set of divisors on $\mathbb{P}^{1}(K)$.
Definition D.3.
For a rational function $f\colon\mathbb{P}^{1}(K)\to\mathbb{P}^{1}(K)$, $\alpha\in\mathbb{P}^{1}(K)$ and a subset $T\subset\mathbb{P}^{1}(K)$, we define $\operatorname{div}(f=\alpha\mid T)\in\operatorname{Div}\mathbb{P}^{1}(K)$ as
@@
where $n_{P}$ is the multiplicity of $f=\alpha$ at $P$. We simply write
$\operatorname{div}(f=\alpha)=\operatorname{div}(f=\alpha\mid\mathbb{P}^{1}(K)).$
@@
Let $f\colon\mathbb{P}^{1}(K)\to\mathbb{P}^{1}(K)$ be a rational function and let $(T_{\lambda})_{\lambda\in\Lambda}$ be a family of disjoint subsets of $\mathbb{P}^{1}(K)$. Then, we have
@@
For a divisor $D=\sum_{P\in\mathbb{P}^{1}(K)}n_{P}P$, we define its degree and support as
$\deg D=\sum_{P\in\mathbb{P}^{1}(K)}n_{P},\quad\operatorname{supp}D=\{P\in\mathbb{P}^{1}(K)\mid n_{P}\neq 0\}.$
@@
For a rational function $f\colon\mathbb{P}^{1}(K)\to\mathbb{P}^{1}(K)$ and an arbitrary value $\alpha\in\mathbb{P}^{1}(K)$, the following equality holds:
@@
where the right-hand side is the degree of $f$ as a rational function. In particular, the degree of a rational function coincides with the number of preimages counted with multiplicity.